OpenStudy (anonymous):

Help this mom please! :) The sum 51 5/6 + (- 37 1/12), is equal to what? Thank u! :) Mom

OpenStudy (jdoe0001):

\(\bf 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\\ \quad \\ \quad \\ {\color{blue}{ 51\cfrac{5}{6}\implies \cfrac{51\cdot 6+5}{6}\implies \cfrac{311}{6}\\ \quad \\ -37\cfrac{1}{12}\implies \cfrac{-37\cdot 12+1}{12}\implies \cfrac{-443}{12}\\ \quad \\}} 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\implies \cfrac{311}{6}+\left(\cfrac{-443}{12}\right)\implies \cfrac{311}{6}-\cfrac{443}{12}\\ \quad \\ \implies \cfrac{622+443}{12}\implies \cfrac{355}{4}\)

OpenStudy (anonymous):

thats what i got, but thought it was wrong! thank so very much :)

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

ohhh ll wait a sec

OpenStudy (jdoe0001):

hold the mayo.. .I did make a mistake.... I use a +, rather than -

OpenStudy (anonymous):

oops forgot to show you the answers to pick from !

OpenStudy (anonymous):

14 3/4, -14 3/4,

OpenStudy (anonymous):

OR...

OpenStudy (jdoe0001):

\(\bf 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\\ \quad \\ \quad \\ {\color{blue}{ 51\cfrac{5}{6}\implies \cfrac{51\cdot 6+5}{6}\implies \cfrac{311}{6}\\ \quad \\ -37\cfrac{1}{12}\implies \cfrac{-37\cdot 12+1}{12}\implies \cfrac{-443}{12}\\ \quad \\}} 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\implies \cfrac{311}{6}+\left(\cfrac{-443}{12}\right)\implies \cfrac{311}{6}-\cfrac{443}{12}\\ \quad \\ \implies \cfrac{622-443}{12}\implies \cfrac{179}{12}\)

OpenStudy (anonymous):

88 3/4, OR 88 11/12

OpenStudy (anonymous):

i am confused

OpenStudy (jdoe0001):

hmmm .... well, on the previous one... I did make a mistake... I use the +... I added the numerator, you do have a minus.... .so ... I sorta made a quick typo and added by mistake

OpenStudy (anonymous):

I dont see any of the possible answers in your equations., or mine...

OpenStudy (jdoe0001):

\(\bf \cfrac{a}{b}+\left(-\cfrac{c}{d}\right)\implies \cfrac{a}{b}-\cfrac{c}{d}\)

OpenStudy (jdoe0001):

ol.... lemme recheck myself

OpenStudy (jdoe0001):

hmm I see my another mistake.... gimme a sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\\ \quad \\ \quad \\ 51\cfrac{5}{6}\implies \cfrac{51\cdot 6+5}{6}\implies {\color{blue}{ \cfrac{311}{6}}}\\ \quad \\ 37\cfrac{1}{12}\implies \cfrac{37\cdot 12+1}{12}\implies {\color{blue}{ \cfrac{445}{12}}}\\ \quad \\ 51\cfrac{5}{6}+\left(-37\cfrac{1}{12}\right)\implies \cfrac{311}{6}+\left(-\cfrac{445}{12}\right)\implies \cfrac{311}{6}-\cfrac{445}{12}\\ \quad \\ \implies \cfrac{622-445}{12}\implies \cfrac{177}{12}\)

OpenStudy (anonymous):

but that answer is not one of the answers to pick from

OpenStudy (anonymous):

i posted the 4 answers above

OpenStudy (jdoe0001):

\(\bf \cfrac{177}{12}\implies {\color{red}{ 14\cfrac{3}{4}}}\implies \cfrac{14\cdot 4+3}{4}\)

OpenStudy (anonymous):

thank u!!

OpenStudy (jdoe0001):

well... ahemm I shoulld say... \(\bf \cfrac{177}{12}\implies \cfrac{59}{4}\implies {\color{red}{ 14\cfrac{3}{4}}}\implies \cfrac{14\cdot 4+3}{4}\)

OpenStudy (jdoe0001):

it simplifies to to \(\bf \cfrac{59}{4}\) and you can write as a mixed fraction of \(\bf 14\cfrac{3}{4}\)

OpenStudy (jdoe0001):

yw